# On the (In)Approximability of the Monitoring Edge Geodetic Set Problem

### v1

@@ 3 Inapproximability of MEG-Set
All polynomial-time approximation algorithms for MEG-Set must have an approximation factor of $\Omega(\log n)$, unless $\mathsf{P}=\mathsf{NP}$. Moreover, assuming $\mathsf{NP}\subseteq\mathsf{DTIME}(n^{O(\log\log n)})$, the approximation factor must be at least $(\frac{1}{2}-\varepsilon)\ln n$ for any constant $\varepsilon>0$.
4 $\mathsf{NP}$-Hardness of minimum MEG-Set on graphs with apex $1$
In this section, we show a reduction from the minimum Planar Dominating Set problem on graphs with girth at least seven (which is known to be $\mathsf{NP}$-Hard [22]) to the minimum MEG-Set on graphs with apex $1$ problem.
Let $G=(V,E)$ be a planar input graph with $|V(G)|\geq 2$ and with girth $g(G)\geq 7$. We construct an instance of MEG-Set on graphs with apex $1$, namely $H$, starting from $G$ and augmenting it as follows: for each vertex $v\in V(G)$, we add 2 vertices $v^{\prime}$ and $v^{\prime\prime}$ to $G$, then connect $v$ to $v^{\prime}$ and $v^{\prime}$ to $v^{\prime\prime}$. We denote the set containing all the vertices $v^{\prime}$ as $L^{\prime}$ and the set containing all the vertices $v^{\prime\prime}$ as $L^{\prime\prime}$. Notice that:
  1. i)
  2. ii)
The basic idea is to monitor each edge that is incident to a vertex $v\in V(G)$ that belongs to a selected MEG-Set in $H$. We denote the set of edges $(v,v^{\prime})$ as $E_{L^{\prime}}$ and the set of edges $(v^{\prime},v^{\prime\prime})$ as $E_{L^{\prime\prime}}$. Note that at this point of the construction, all edges are monitored by at least one pair of leaves in $L^{\prime\prime}$. To avoid this, we shortcut all the shortest paths from one vertex in $L^{\prime\prime}$ to another by inserting two vertices $v_{*}$ and $v_{*}^{\prime}$ and connecting them together. Then, we connect $v_{*}$ to all vertices $v^{\prime}\in L^{\prime}$. We denote the set containing edges $(v_{*},v^{\prime})$ with $v^{\prime}\in L^{\prime}$ as $E_{*}$. The vertex $v_{*}$ is used to create shortcuts between each shortest path from a vertex in $L^{\prime}\cup L^{\prime\prime}$ to a vertex in $L^{\prime}\cup L^{\prime\prime}$. Notice that, since $v_{*}^{\prime}$ is a leaf, it belongs to every MEG-Set of $G$ due to Lemma 2.1. The vertex $v_{*}^{\prime}$ is used to monitor the edges in $E_{*}$. Formally, let $H=(V(G)\cup L^{\prime}\cup L^{\prime\prime}\cup\{v_{*},v_{*}^{\prime}\},E(G)\cup E_{L^{\prime}}\cup E_{L^{\prime\prime}}\cup E_{*}\cup\{\;\{v_{*},v_{*}^{\prime}\}\;\})$, where:
  • $L^{\prime}=\{v^{\prime}:v\in V(G)\}$;
  • $L^{\prime\prime}=\{v^{\prime\prime}:v\in V(G)\}$;
  • $E_{L^{\prime}}=\{\{v,v^{\prime}\}:v\in V(G)\}$;
  • $E_{L^{\prime\prime}}=\{\{v^{\prime\prime},v^{\prime}\}:v\in V(G)\}$;
  • $E_{*}=\{\{v_{*},v^{\prime}\}:v\in V(G)\}$.
$a$$b$$c$$d$$a^{\prime}$$b^{\prime}$$c^{\prime}$$d^{\prime}$$a^{\prime\prime}$$b^{\prime\prime}$$c^{\prime\prime}$$d^{\prime\prime}$$v_{*}$$v_{*}^{\prime}$
Figure 2: An example of the construction of the graph $H$ from the input graph $G=(\{a,b,c,d\},\{\;\{a,b\},\{b,c\},\{c,d\},\{d,a\}\;\})$ (shown in black).
We prove the $\mathsf{NP}$-Completeness of minimum MEG-Set on graphs with apex $1$ via two lemmas. The first lemma proves that if there is a MEG-Set of the graph $H$ having size at most $k+|V(G)|+1$, there exists a Dominating Set of $G$ having size at most $k$. Conversely, the second lemma proves that if there is a Dominating Set of the graph $G$ having size at most $k$, there exists a MEG-Set of $H$ having size $k+|V(G)|+1$.
Lemma 4.1.
Proof 4.2.
  • Case 3: $e\in E_{L^{\prime}}$. Let $e=(u,u^{\prime})$ where $u$ is the unique endpoint of $e$ in $V(G)$. Since $D$ dominates all the vertices in $V(G)$, either $u\in D$, and we let $x=u$, or $u\not\in D$ and there exists some vertex $x\in D\cap N(u)$. There is a single shortest path from $u^{\prime\prime}$ to $x$ and it traverses the edge $(u,u^{\prime})$.
Lemma 4.3.
Proof 4.4.
Theorem 4.5.
Proof 4.6.
For each planar graph $G$ with girth at least seven, it is possible to construct in polynomial time a graph $H$ with apex $1$, as discussed in this section. From Lemmas 4.1 and 4.3, there exists a Dominating Set $D$ of $G$ such that $|D|\leq k$ if and only if there exists a MEG-Set $M$ of $H$ such that $|M|\leq k+|V(G)|+1$. Since the Dominating Set problem is $\mathsf{NP}$-Hard even for planar graphs with girth at least seven, this implies that the MEG-Set decision problem is $\mathsf{NP}$-Hard on graphs with apex $1$.
@@ 5.1 A simple approximation algorithm for GMEG-Set
Let $Z$ be the set of all vertices of cost 00. Observing that it is possible to check in polynomial time whether an edge of $E^{\prime}$ is monitored by $Z$, we construct an instance of Set Cover in which the items are exactly the edges in $E^{\prime}$ that are not monitored by $Z$. The instance contains a set $S_{p}$ for each (unordered) pair $p=\{u,v\}$ of distinct vertices $u,v$ in $V\setminus Z$. The set $S_{p}$ contains all items $e$ that are monitored by $Z\cup p$ (notice that an item in $S_{p}$ is not necessarily monitored by $p$).
@@
Lemma 5.2 implies the existence of a polynomial-time $O(\sqrt{n\log n})$-approximation algorithm for MEG-Set because $c(V)=n$.
@@ 5.2.1 Some structural properties
Given a cluster $C$, we define the projection of $C$ as the following instance of GMEG-Set. The instance’s graph $H$ has a vertex set consisting of all vertices in $C\cup\delta C$, where vertices in $C$ retain their original cost in $G$, while vertices in $\delta C$ have a cost of 00. The edge set of $H$ consists of all edges in $G[C\cup\delta C]$ with their original weight in $G$, plus some additional projection edges. More precisely, for each pair of distinct vertices $u,v\in\delta C$ with at least one $C$-bypass between them, we add a projection edge $e=(u,v)$ of weight $d_{G}(u,v)$, and we associate $e$ with an arbitrarily selected $C$-bypass $b_{e}$ between $u$ and $v$. We refer to $H$ as the projection graph of $C$, and we use $E_{C}$ to denote the set of all non-projection edges of $H$. The edges of $H$ to be monitored are all those in $E(H)\cap E^{\prime}$. See Figure 3 for an example.
Figure 3: (a) The graph $G$ of an instance of GMEG-Set and a cluster $C$. Vertices in the boundary $\delta C$ are shown in bold. Three $C$-bypasses are highlighted in red, blue, and green, respectively. All edge weights are unitary and are not shown. (b) The graph of the projection $H$ of $C$. There are three projection edges, highlighted in red, blue, and green, corresponding to the $C$-bypass shown with the same color in (a). The number next to a projection edge represents its weight, while non-projection edges have unit weight (not shown). Gray vertices form a GMEG-Set of $H$. The vertices in the boundary are shown in bold and have a cost of 00. Notice how the gray vertices also monitor all non-projection edges in $G$.
@@
We start by proving (a). Let $\pi$ be a shortest path between $u$ and $v$ in $H$. By Lemma 5.3 (a), there exists a walk $\pi^{\prime}$ between $u$ and $v$ in $G$ such that $w_{H}(\pi)=w_{G}(\pi^{\prime})$ and $E(\pi)\cap E_{C}=E(\pi^{\prime})\cap E_{C}$. We now argue that $\pi^{\prime}$ is actually a shortest path in $G$. Suppose, towards a contradiction, that there exists some path $\pi^{\prime\prime}$ between $u$ and $v$ in $G$ such that $w_{G}(\pi^{\prime\prime})<w_{G}(\pi^{\prime})$. By Lemma 5.3 (b), this implies the existence of a walk $\pi^{\prime\prime\prime}$ between $u$ and $v$ in $H$ such that $w_{H}(\pi^{\prime\prime\prime})=w_{G}(\pi^{\prime\prime})$, yielding the contradiction $w_{H}(\pi)=w_{G}(\pi^{\prime})>w_{G}(\pi^{\prime\prime})=w_{H}(\pi^{\prime\prime\prime})$.
We now prove (b). Let $\pi$ be a shortest path between $u$ and $v$ in $G$. By Lemma 5.3 (b), there exists a walk $\pi^{\prime}$ between $u$ and $v$ in $H$ such that $w_{G}(\pi)=w_{H}(\pi^{\prime})$ and $E(\pi)\cap E_{C}=E(\pi^{\prime})\cap E_{C}$. We now argue that $\pi^{\prime}$ is actually a shortest path in $H$. Suppose, towards a contradiction, that there exists some path $\pi^{\prime\prime}$ between $u$ and $v$ in $H$ such that $w_{H}(\pi^{\prime\prime})<w_{H}(\pi^{\prime})$. By Lemma 5.3 (a), this implies the existence of a walk $\pi^{\prime\prime\prime}$ between $u$ and $v$ in $G$ such that $w_{G}(\pi^{\prime\prime\prime})=w_{H}(\pi^{\prime\prime})$, yielding the contradiction $w_{G}(\pi)=w_{H}(\pi^{\prime})>w_{H}(\pi^{\prime\prime})=w_{G}(\pi^{\prime\prime\prime})$.
@@
Assume that $e$ is monitored by the pair $u,v$ in $G$, and let $\pi$ be a shortest path between $u$ and $v$ in $G$. By Corollary 5.5 (b), there exists a shortest path $\pi^{\prime}$ between $u$ and $v$ in $H$ that traverses $e$. We now argue that all shortest paths between $u$ and $v$ in $H$ contain $e$. Indeed, if $H$ contained a shortest path $\pi^{\prime\prime}$ between $u$ and $v$ avoiding $e$, then Corollary 5.5 (a) would imply the existence of a shortest path between $u$ and $v$ avoiding $e$ in $G$, i.e., $u,v$ would not monitor $e$ in $G$.
Assume now that $e$ is monitored by the pair $u,v$ in $H$, and let $\pi$ be a shortest path between $u$ and $v$ in $H$. By Corollary 5.5 (a), there exists a shortest path $\pi^{\prime}$ between $u$ and $v$ in $G$ that traverses $e$. We now argue that all shortest paths between $u$ and $v$ in $G$ contain $e$. Indeed, if $G$ contained a shortest path $\pi^{\prime\prime}$ between $u$ and $v$ avoiding $e$, then Corollary 5.5 (b) would imply the existence of a shortest path between $u$ and $v$ avoiding $e$ in $H$, i.e., $u,v$ would not monitor $e$ in $H$.
@@
Since the subpath of $\pi$ from $u$ to $v$ contains edge $e$, the pair $u,v$ monitors $e$ in $G$. Hence, by Corollary 5.7, $u,v$ monitors $e$ in the projection graph of $C$.
@@
Let $M$ be a GMEG-Set for the projection of $C$, and let $H$ be the corresponding graph. We consider a generic edge $e\in E_{C}\cap E^{\prime}$ and we prove that $e$ is also monitored by $M$ in $G$. Let $u,v\in M$ be a pair of vertices that monitors $e$ in $H$. By Corollary 5.7, the pair $u,v$ monitors $e$ in $G$.
@@ 5.2.2 Algorithm description
[…]
@@
Let $G^{\prime}$ be the induced subgraph of $G$ that contains all (and only) the vertices having non-zero cost. We build a rooted tree $T$ in which each vertex $u$ of $T$ is associated with some subgraph $G_{u}$ of $G^{\prime}$ and with an instance of GMEG-Set $H_{u}$. The root $r$ of $T$ is associated with $G_{r}=G^{\prime}$ and with the projection $H_{r}$ of $V(G_{r})$ (w.r.t. the input instance of GMEG-Set). A generic vertex $u$ of $T$ is a leaf if $H_{u}$ admits a solution of cost 00, or if $G_{u}$ is $\alpha$-non-separable. Otherwise, $u$ has exactly two children $a,b$ in $T$ corresponding to the subgraphs $G_{a}=G_{u}[A]$ and $G_{b}=G_{u}[B]$ induced by the sets $A,B$ of an $\alpha$-balanced vertex-separator $S_{u},A,B$ of $G_{u}$. The instances of GMEG-Set $H_{a}$ and $H_{b}$ are the projections of $V(G_{a})$ and $V(G_{b})$ w.r.t. the instances $H_{u}$, respectively. See Figure 4 for an example.
Figure 4: The tree $T$ representing a possible hierarchical decomposition of the graph shown in the root vertex for the parameters $\beta_{0}=1$, $\beta=\frac{1}{2}$, $\alpha=\frac{1}{4}$. Vertices in a separator are shown in gray in the corresponding cluster. Black vertices in a leaf cluster $u$ of $T$ form an optimal solution to instance $H_{u}$ (not shown) when combined with the vertices of cost 00 resulting from the projection. Observe that the instance corresponding to the right child of the root admits a solution of cost 00. Gray and black vertices in an internal cluster $u$ form a feasible GMEG-Set of the corresponding instance $H_{u}$ (not shown) when combined with the vertices of cost 00.
@@
In the second phase of our algorithm, we visit the tree in a bottom-up fashion and, for each examined vertex $u$, we compute a $\rho(h_{u},n_{u})$-approximate solution for the projection $H_{u}$ of $V(G_{u})$. Here, $\rho(h_{u},n_{u})$ is a function that depends on the number $n_{u}=|V(G_{u})|$ of vertices in $G_{u}$ and, possibly, on the height $h_{u}$ of $u$ in $T$ [^6] The height of vertex $u$ in $T$ is the height of the subtree of $T$ rooted in $u$, i.e., the maximum number of edges in the longest path from $u$ to a (not necessarily proper) descendant of $u$. Observe that all vertices in $G_{u}$ have cost $1$ w.r.t. the cost function $c_{u}$ of $H_{u}$, i.e., $n_{u}=c_{u}(V(G_{u}))$.
In detail, if $u$ is a leaf in $T$, then either there exists an optimal GMEG-Set of $H_{u}$ with cost 00, or $|V(G_{u})|=O(1)$, and we can find an optimal GMEG-Set of $H_{u}$ in constant time by brute force (possibly both). Regardless of the case, we have an optimal solution for $H_{u}$, i.e., a $\rho(0,n_{u})$-approximation for $\rho(0,n_{u})=1$.
When the examined vertex $u$ of $T$ is a generic internal vertex at height $h$ having $a$ and $b$ as children, we compute a GMEG-Set $M_{u}$ for $H_{u}$ by choosing the set of smallest cost between (i) a GMEG-Set $M^{\prime}_{u}$ obtained as the union of a $\rho(h_{u}-1,n_{a})$-approximate GMEG-Set $M_{a}$ for $H_{a}$ with a $\rho(h_{u}-1,n_{b})$-approximate GMEG-Set $M_{b}$ for $H_{b}$, and (ii) a $(2c_{u}(M^{*})\ln|V(G_{u})|)$-approximate solution computed using the algorithm of Section 5.1 (see Lemma 5.1), where $M^{*}$ denotes an optimal solution for $H_{u}$ and $c_{u}$ is the cost function of $H_{u}$. The next lemma provides an upper bound on the cost of $c_{u}(M^{\prime}_{u})$.
@@
Let $M^{*}_{a}=(M^{*}\cap V(G_{a}))\cup\delta V(G_{a})$ (resp. $M^{*}_{b}=(M^{*}\cap V(G_{b}))\cup\delta V(G_{b})$), where the boundary is w.r.t. the instance $H_{u}$. By Lemma 5.9, $M^{*}_{a}$ and $M^{*}_{b}$ are GMEG-Set for $H_{a}$ and $H_{b}$, respectively. Then, $c_{a}(M^{*}_{a})=c_{a}(M^{*}\cap V(G_{a}))+c_{a}(\delta V(G_{a}))=c_{a}(M^{*}\cap V(G_{a}))=c_{u}(M^{*}\cap V(G_{a}))$, and a similar derivation shows that $c_{b}(M^{*}_{b})=c_{u}(M^{*}\cap V(G_{b}))$.
@@
Let $E^{\prime}_{u}$ be the set of edges to be monitored in $H_{u}$. By Lemma 5.11, $M_{a}$ monitors all edges of $E^{\prime}_{u}$ in $G_{u}[V(G_{a})\cup\delta V(G_{a})]$, and $M_{b}$ monitors all edges of $E^{\prime}_{u}$ in $G_{u}[V(G_{b})\cup\delta V(G_{b})]$; therefore, $M_{a}\cup M_{b}$ is a GMEG-Set for $H_{u}$. Since $M_{a}\cup M_{b}\subseteq M^{\prime}_{u}$, we have that $M^{\prime}_{u}$ is a GMEG-Set of $H_{u}$ of cost:
@@
The proof is by induction on $h_{u}$. The claim is trivially true for $h_{u}=0$ since $\rho(0,n_{u})=1\leq\sqrt{2\ln 2}\leq\frac{\sqrt{n_{u}\beta_{0}\ln n}}{1-(1-\alpha)^{\beta/2}}$. Therefore, we assume that the claim holds for $h_{u}-1\geq 0$ and we prove that it holds for $h_{u}$. Consider a vertex $u$ at height $h_{u}$, and let $a$ and $b$ be its two children, and let $S_{u}$ be the corresponding separator of $G_{u}$. Let $M^{*}$ be an optimal GMEG-Set for $H_{u}$. By Lemma 5.14, we have:
@@
The proof is by induction on $h$. The claim is trivially true for $h=0$ since $\rho(0,n_{u})=1\leq\sqrt{2\ln 2}\leq(h_{u}+1)\sqrt{n_{u}\beta_{0}\ln n}$. Therefore, we assume that the claim holds for $h_{u}-1\geq 0$ and we prove that it holds for $h_{u}$. Consider a vertex $u$ at height $h_{u}$, let $a$ and $b$ be its two children, and let $S_{u}$ be the corresponding separator of $G_{u}$. Let $M^{*}$ be an optimal solution for $H_{u}$. By Lemma 5.14, we have:

### v3

@@ 4 Inapproximability of MEG-Set
All polynomial-time approximation algorithms for MEG-Set must have an approximation factor of $\Omega(\log n)$, unless $\mathsf{P}=\mathsf{NP}$. Moreover, unless $\mathsf{NP}\subseteq\mathsf{DTIME}(n^{O(\log\log n)})$, the approximation factor must be at least $(\frac{1}{2}-\varepsilon)\ln n$ for any constant $\varepsilon>0$.
@@ 5.1 A simple approximation algorithm for GMEG-Set
Let $Z$ be the set of all vertices of cost 0. Observing that it is possible to check in polynomial time whether an edge of $E^{\prime}$ is monitored by $Z$, we construct an instance of Set Cover in which the items are exactly the edges in $E^{\prime}$ that are not monitored by $Z$. The instance contains a set $S_{p}$ for each (unordered) pair $p=\{u,v\}$ of distinct vertices $u,v$ in $V\setminus Z$. The set $S_{p}$ contains all items $e$ that are monitored by $Z\cup p$ (notice that an item in $S_{p}$ is not necessarily monitored by $p$).
@@
Lemma˜5.2 implies the existence of a polynomial-time $O(\sqrt{n\log n})$-approximation algorithm for MEG-Set because $c(V)=n$.
@@ 5.2.1 Some structural properties
Given a cluster $C$, we define the projection of $C$ as the following instance of GMEG-Set. The instance’s graph $H$ has a vertex set consisting of all vertices in $C\cup\delta C$, where vertices in $C$ retain their original cost in $G$, while vertices in $\delta C$ have a cost of 0. The edge set of $H$ consists of all edges in $G[C\cup\delta C]$ with their original weight in $G$, plus some additional projection edges. More precisely, for each pair of distinct vertices $u,v\in\delta C$ with at least one $C$-bypass between them, we add a projection edge $e=(u,v)$ of weight $d_{G}(u,v)$, and we associate $e$ with an arbitrarily selected $C$-bypass $b_{e}$ between $u$ and $v$. We refer to $H$ as the projection graph of $C$, and we use $E_{C}$ to denote the set of all non-projection edges of $H$. The edges of $H$ to be monitored are all those in $E(H)\cap E^{\prime}$. See Figure˜3 for an example.
Figure 3: (a) The graph $G$ of an instance of GMEG-Set and a cluster $C$. Vertices in the boundary $\delta C$ are shown in bold. Three $C$-bypasses are highlighted in red, blue, and green, respectively. All edge weights are unitary and are not shown. (b) The graph of the projection $H$ of $C$. There are three projection edges, highlighted in red, blue, and green, corresponding to the $C$-bypass shown with the same color in (a). The number next to a projection edge represents its weight, while non-projection edges have unit weight (not shown). Gray vertices form a GMEG-Set of $H$. The vertices in the boundary are shown in bold and have a cost of 0. Notice how the gray vertices also monitor all non-projection edges in $G$.
@@
We start by proving (a). Let $\pi$ be a shortest path between $u$ and $v$ in $H$. By Lemma˜5.3 (a), there exists a walk $\pi^{\prime}$ between $u$ and $v$ in $G$ such that $w_{H}(\pi)=w_{G}(\pi^{\prime})$ and $E(\pi)\cap E_{C}=E(\pi^{\prime})\cap E_{C}$. We now argue that $\pi^{\prime}$ is actually a shortest path in $G$. Suppose, towards a contradiction, that there exists some path $\pi^{\prime\prime}$ between $u$ and $v$ in $G$ such that $w_{G}(\pi^{\prime\prime})<w_{G}(\pi^{\prime})$. By Lemma˜5.3 (b), this implies the existence of a walk $\pi^{\prime\prime\prime}$ between $u$ and $v$ in $H$ such that $w_{H}(\pi^{\prime\prime\prime})=w_{G}(\pi^{\prime\prime})$, yielding the contradiction $w_{H}(\pi)=w_{G}(\pi^{\prime})>w_{G}(\pi^{\prime\prime})=w_{H}(\pi^{\prime\prime\prime})$.
We now prove (b). Let $\pi$ be a shortest path between $u$ and $v$ in $G$. By Lemma˜5.3 (b), there exists a walk $\pi^{\prime}$ between $u$ and $v$ in $H$ such that $w_{G}(\pi)=w_{H}(\pi^{\prime})$ and $E(\pi)\cap E_{C}=E(\pi^{\prime})\cap E_{C}$. We now argue that $\pi^{\prime}$ is actually a shortest path in $H$. Suppose, towards a contradiction, that there exists some path $\pi^{\prime\prime}$ between $u$ and $v$ in $H$ such that $w_{H}(\pi^{\prime\prime})<w_{H}(\pi^{\prime})$. By Lemma˜5.3 (a), this implies the existence of a walk $\pi^{\prime\prime\prime}$ between $u$ and $v$ in $G$ such that $w_{G}(\pi^{\prime\prime\prime})=w_{H}(\pi^{\prime\prime})$, yielding the contradiction $w_{G}(\pi)=w_{H}(\pi^{\prime})>w_{H}(\pi^{\prime\prime})=w_{G}(\pi^{\prime\prime\prime})$.
@@
Assume that $e$ is monitored by the pair $u,v$ in $G$, and let $\pi$ be a shortest path between $u$ and $v$ in $G$. By Corollary˜5.5 (b), there exists a shortest path $\pi^{\prime}$ between $u$ and $v$ in $H$ that traverses $e$. We now argue that all shortest paths between $u$ and $v$ in $H$ contain $e$. Indeed, if $H$ contained a shortest path $\pi^{\prime\prime}$ between $u$ and $v$ avoiding $e$, then Corollary˜5.5 (a) would imply the existence of a shortest path between $u$ and $v$ avoiding $e$ in $G$, i.e., $u,v$ would not monitor $e$ in $G$.
Assume now that $e$ is monitored by the pair $u,v$ in $H$, and let $\pi$ be a shortest path between $u$ and $v$ in $H$. By Corollary˜5.5 (a), there exists a shortest path $\pi^{\prime}$ between $u$ and $v$ in $G$ that traverses $e$. We now argue that all shortest paths between $u$ and $v$ in $G$ contain $e$. Indeed, if $G$ contained a shortest path $\pi^{\prime\prime}$ between $u$ and $v$ avoiding $e$, then Corollary˜5.5 (b) would imply the existence of a shortest path between $u$ and $v$ avoiding $e$ in $H$, i.e., $u,v$ would not monitor $e$ in $H$.
@@
Since the subpath of $\pi$ from $u$ to $v$ contains edge $e$, the pair $u,v$ monitors $e$ in $G$. Hence, by Corollary˜5.7, $u,v$ monitors $e$ in the projection graph of $C$.
@@
Let $M$ be a GMEG-Set for the projection of $C$, and let $H$ be the corresponding graph. We consider a generic edge $e\in E_{C}\cap E^{\prime}$ and we prove that $e$ is also monitored by $M$ in $G$. Let $u,v\in M$ be a pair of vertices that monitors $e$ in $H$. By Corollary˜5.7, the pair $u,v$ monitors $e$ in $G$.
@@ 5.2.2 Algorithm description
We consider a class $\mathcal{G}$ with efficiently computable $\alpha$-balanced separators of size $\beta_{0}|V(G)|^{\beta}$ for some constants $\beta_{0}\geq 1$ and $\beta\geq 0$, and we restrict our attention (w.l.o.g.) to instances of GMEG-Set whose input graph $G=(V,E)$ belongs to $\mathcal{G}$ and satisfies $c(V)\geq 2$. [^5] Otherwise, when $c(V)\leq 1$, an optimal solution consists either of all vertices of cost 0, or of all vertices of cost 0 plus the sole vertex of cost $1$. Our algorithm works in two phases.
@@
Let $G^{\prime}$ be the induced subgraph of $G$ that contains all (and only) the vertices having non-zero cost. We build a rooted tree $T$ in which each vertex $u$ of $T$ is associated with some subgraph $G_{u}$ of $G^{\prime}$ and with an instance of GMEG-Set $H_{u}$. The root $r$ of $T$ is associated with $G_{r}=G^{\prime}$ and with the projection $H_{r}$ of $V(G_{r})$ (w.r.t. the input instance of GMEG-Set). A generic vertex $u$ of $T$ is a leaf if $H_{u}$ admits a solution of cost 0, or if $G_{u}$ is $\alpha$-non-separable. Otherwise, $u$ has exactly two children $a,b$ in $T$ corresponding to the subgraphs $G_{a}=G_{u}[A]$ and $G_{b}=G_{u}[B]$ induced by the sets $A,B$ of an $\alpha$-balanced vertex-separator $S_{u},A,B$ of $G_{u}$. The instances of GMEG-Set $H_{a}$ and $H_{b}$ are the projections of $V(G_{a})$ and $V(G_{b})$ w.r.t. the instances $H_{u}$, respectively. See Figure˜4 for an example.
Figure 4: The tree $T$ representing a possible hierarchical decomposition of the graph shown in the root vertex for the parameters $\beta_{0}=1$, $\beta=\frac{1}{2}$, $\alpha=\frac{1}{4}$. Vertices in a separator are shown in gray in the corresponding cluster. Black vertices in a leaf cluster $u$ of $T$ form an optimal solution to instance $H_{u}$ (not shown) when combined with the vertices of cost 0 resulting from the projection. Observe that the instance corresponding to the right child of the root admits a solution of cost 0. Gray and black vertices in an internal cluster $u$ form a feasible GMEG-Set of the corresponding instance $H_{u}$ (not shown) when combined with the vertices of cost 0.
@@
In the second phase of our algorithm, we visit the tree in a bottom-up fashion and, for each examined vertex $u$, we compute a $\rho(h_{u},n_{u})$-approximate solution for the projection $H_{u}$ of $V(G_{u})$. Here, $\rho(h_{u},n_{u})$ is a function that depends on the number $n_{u}=|V(G_{u})|$ of vertices in $G_{u}$ and, possibly, on the height $h_{u}$ of $u$ in TT. [^6] The height of vertex $u$ in $T$ is the height of the subtree of $T$ rooted in $u$, i.e., the maximum number of edges in the longest path from $u$ to a (not necessarily proper) descendant of $u$. Observe that all vertices in $G_{u}$ have cost $1$ w.r.t. the cost function $c_{u}$ of $H_{u}$, i.e., $n_{u}=c_{u}(V(G_{u}))$.
In detail, if $u$ is a leaf in $T$, then either there exists an optimal GMEG-Set of $H_{u}$ with cost 0, or $|V(G_{u})|=O(1)$, and we can find an optimal GMEG-Set of $H_{u}$ in constant time by brute force (possibly both). Regardless of the case, we have an optimal solution for $H_{u}$, i.e., a $\rho(0,n_{u})$-approximation for $\rho(0,n_{u})=1$.
When the examined vertex $u$ of $T$ is a generic internal vertex at height $h$ having $a$ and $b$ as children, we compute a GMEG-Set $M_{u}$ for $H_{u}$ by choosing the set of smallest cost between (i) a GMEG-Set $M^{\prime}_{u}$ obtained as the union of a $\rho(h_{u}-1,n_{a})$-approximate GMEG-Set $M_{a}$ for $H_{a}$ with a $\rho(h_{u}-1,n_{b})$-approximate GMEG-Set $M_{b}$ for $H_{b}$, and (ii) a $(2c_{u}(M^{*})\ln|V(G_{u})|)$-approximate solution computed using the algorithm of Section˜5.1 (see Lemma˜5.1), where $M^{*}$ denotes an optimal solution for $H_{u}$ and $c_{u}$ is the cost function of $H_{u}$. The next lemma provides an upper bound on the cost of $c_{u}(M^{\prime}_{u})$.
@@
Let $M^{*}_{a}=(M^{*}\cap V(G_{a}))\cup\delta V(G_{a})$ (resp. $M^{*}_{b}=(M^{*}\cap V(G_{b}))\cup\delta V(G_{b})$), where the boundary is w.r.t. the instance $H_{u}$. By Lemma˜5.9, $M^{*}_{a}$ and $M^{*}_{b}$ are GMEG-Set for $H_{a}$ and $H_{b}$, respectively. Then, $c_{a}(M^{*}_{a})=c_{a}(M^{*}\cap V(G_{a}))+c_{a}(\delta V(G_{a}))=c_{a}(M^{*}\cap V(G_{a}))=c_{u}(M^{*}\cap V(G_{a}))$, and a similar derivation shows that $c_{b}(M^{*}_{b})=c_{u}(M^{*}\cap V(G_{b}))$.
@@
Let $E^{\prime}_{u}$ be the set of edges to be monitored in $H_{u}$. By Lemma˜5.11, $M_{a}$ monitors all edges of $E^{\prime}_{u}$ in $G_{u}[V(G_{a})\cup\delta V(G_{a})]$, and $M_{b}$ monitors all edges of $E^{\prime}_{u}$ in $G_{u}[V(G_{b})\cup\delta V(G_{b})]$; therefore, $M_{a}\cup M_{b}$ is a GMEG-Set for $H_{u}$. Since $M_{a}\cup M_{b}\subseteq M^{\prime}_{u}$, we have that $M^{\prime}_{u}$ is a GMEG-Set of $H_{u}$ of cost:
@@
The proof is by induction on $h_{u}$. The claim is trivially true for $h_{u}=0$ since $\rho(0,n_{u})=1\leq\sqrt{2\ln 2}\leq\frac{\sqrt{n_{u}\beta_{0}\ln n}}{1-(1-\alpha)^{\beta/2}}$. Therefore, we assume that the claim holds for $h_{u}-1\geq 0$ and we prove that it holds for $h_{u}$. Consider a vertex $u$ at height $h_{u}$, and let $a$ and $b$ be its two children, and let $S_{u}$ be the corresponding separator of $G_{u}$. Let $M^{*}$ be an optimal GMEG-Set for $H_{u}$. By Lemma˜5.14, we have:
@@
The proof is by induction on $h$. The claim is trivially true for $h=0$ since $\rho(0,n_{u})=1\leq\sqrt{2\ln 2}\leq(h_{u}+1)\sqrt{n_{u}\beta_{0}\ln n}$. Therefore, we assume that the claim holds for $h_{u}-1\geq 0$ and we prove that it holds for $h_{u}$. Consider a vertex $u$ at height $h_{u}$, let $a$ and $b$ be its two children, and let $S_{u}$ be the corresponding separator of $G_{u}$. Let $M^{*}$ be an optimal solution for $H_{u}$. By Lemma˜5.14, we have: